# Growing a Twig to Accelerate Large Vision-Language Models

### v1

@@ Abstract
Large vision-language models (VLMs) have demonstrated remarkable capabilities in open-world multimodal understanding, yet their high computational overheads pose great challenges for practical deployment. Some recent works have proposed methods to accelerate VLMs by pruning redundant visual tokens guided by the attention maps of VLM’s early layers. Despite the success of these token pruning methods, they still suffer from two major shortcomings: (i) considerable accuracy drop due to insensitive attention signals in early layers, and (ii) limited speedup when generating long responses (e.g., 30 tokens). To address the limitations above, we present TwigVLM—a simple and general architecture by “growing” a lightweight twig upon an early layer of the base VLM. Compared with most existing VLM acceleration methods purely based on visual token pruning, our TwigVLM not only achieves better accuracy retention by employing a twig-guided token pruning (TTP) strategy, but also yields higher generation speed by utilizing a self-speculative decoding (SSD) strategy. Taking LLaVA-1.5-7B as the base VLM, experimental results show that TwigVLM preserves 96% of the original performance after pruning 88.9% of visual tokens and achieves 154% speedup in generating long responses, delivering significantly better performance in terms of both accuracy and speed over the state-of-the-art VLM acceleration methods. Code will be made publicly available.
@@ 1 Introduction
Figure 1: Our TwigVLM overview. (a) Given a base deep VLM, the TwigVLM is obtained by freezing the whole VLM while training a shallow twig block upon an early layer of the VLM. (b) Compared to the VLM acceleration methods based on visual token pruning (e.g., FastV [8]), TwigVLM not only achieves better accuracy retention by employing a twig-guided token pruning (TTP) strategy, but also yields higher generation speed by utilizing a self-speculative decoding (SSD) strategy [12]. Both FastV and TwigVLM take LLaVA-1.5-7B [30] as the base VLM (red dotted lines in the bar graphs) with the same pruning ratio of 88.9%.
The revolution in large language models (LLMs) has fundamentally reshaped the landscape of artificial intelligence [4, 44, 17]. Meanwhile, there has been growing interest in building large vision-language models (VLMs) [39, 68, 29, 1, 9, 53, 25], which can perform general-purpose visual understanding and reasoning and be used in a wide range of tasks such as object grounding [40, 36], document understanding [54, 35], GUI controlling [47, 60], and robotics [3, 22]. Despite the remarkable progress achieved by these large VLMs, they are usually parametric-intensive and computational-heavy. This poses great challenges for using these models in computation-constrained platforms or low-latency scenarios. Therefore, the research on VLM acceleration is of practical and emergent demands.
As the computational cost of the transformer architecture [46] increase quadratically with the length of token sequence, an effective way to accelerate VLMs is to reduce the length of tokens. Moreover, visual tokens usually contain much more redundant information than textual tokens [8], especially in high-resolution images and long videos. Therefore, a series of works on visual token pruning has been proposed [8, 62, 32, 52]. For instance, FastV [8] leverages attention signals from an early layer of the VLM to drop redundant tokens for subsequent layers. Many following works have extended FastV’s idea to progressive token reduction across layers via pruning and merging operations [62, 32], while some others have been proposed to use the attention maps from the visual encoder to guide the visual token pruning before being input into the VLM backbone [52, 61]. Despite the success they have achieved, they still suffer from the following two major shortcomings:
@@
To this end, we present TwigVLM, a simple and general VLM acceleration approach by appending a lightweight twig block upon an early layer of the base VLM (see Figure 1(a)). After efficient post-training of the twig block only, the learned twig block is used in both the prefilling and decoding stages of inference. As shown in Figure 1(b), our TwigVLM not only achieves better accuracy retention by employing a twig-guided token pruning (TTP) strategy, but also yields higher generation speed by utilizing a self-speculative decoding (SSD) strategy [12]. To summarize, the synergistic effect of the TTP and SSD strategies enables holistic inference acceleration of both the prefilling and decoding stages at maximum accuracy retention.
@@ 4.3 Main Results
Results on LLaVA-1.5 and LLaVA-NeXT. As shown in Tables 1 and 2, we take LLaVA-1.5-7B and LLaVA-NeXT-7B as the base VLMs and compare TwigVLM with the state-of-the-art VLM acceleration methods on six commonly-used VLM benchmarks. We can see that TwigVLM consistently and significantly outperforms all its counterparts under different pruning ratios. Notably, TwigVLM achieves near-perfect performance preservation (more than 99.0%) when pruning 66.7% visual tokens on LLaVA-1.5-7B and 77.8% tokens on LLaVA-NeXT-7B, showing the superiority of our TTP strategy over existing pruning strategies. Comparisons on more benchmarks and base VLMs are provided in the supplementary.
@@ 4.4 Ablation Studies
Number of twig layers. Table 4(d) varies the number of twig layers $T$ from 1 to 4 to investigate its effect on TwigVLM. We observe that a single twig layer ($T$=1) leads to the highest speed but the lowest accuracy. Increasing $T$ to 2 or 3 improves accuracy with negligible speed degradation. However, further increasing $T$ to 4 leads to a distinct performance drop in terms of both accuracy and speed. This suggests that $T$=3 achieves an optimal balance in providing more accurate attention signals for pruning (via more twig layers) while still maintaining computational efficiency due to the high token acceptance ratio in SSD.
@@ 5 Related Work
Visual token reduction for VLMs. Unlike acceleration techniques for LLMs, token reduction methods for VLMs mainly address the problem of visual token redundancy to improve inference efficiency [65]. The pioneering work of FastV [8] first studies this redundancy phenomenon, revealing that pruning visual tokens at an early layer of the VLM can achieve acceleration while preserving the majority of the model’s accuracy. Following this paradigm, subsequent studies have introduced adaptive or hierarchical pruning strategies [62, 55, 32, 56, 48, 18] and training-based adoption [7, 50, 64] to further reduce the accuracy drop. Meanwhile, there is another line of studies that performs more aggressive token reduction by pruning or merging visual tokens in the visual encoding stage [2, 41, 52, 61, 45]. However, the accuracy of these pruned models is suboptimal as they are unaware of the textual prompt when pruning visual tokens. Finally, although the above methods effectively accelerate the prefilling stage of VLMs, their speedup in the decoding stage is limited.
@@ Appendix A More Implementation Details
In our experiments, we utilize the LLaVA-665K dataset [30] to train the twig blocks for the LLaVA-1.5-7B and LLaVA-NeXT-7B models, and combine the LLaVA-665K and VideoInstruct-100K datasets [26, 37] for the Video-LLaVA-7B model. The optimization hyper-parameters are detailed in Table 5. All training is performed on a server equipped with 8 NVIDIA A100 GPUs. Under these conditions, the training process is highly efficient, requiring only approximately 10% of the time needed to train the corresponding base VLM, e.g., training the twig block for the LLaVA-1.5-7B model takes about 1 hour, while the training of the original LLaVA-1.5-7B takes about 12 hours.
@@
ablation variant   TokAR (%) RelSpd (%)
Twig block initialization (Table 4c in main text)
(a) random init. 37.7 120.4
(b) VLM layers[$L$-$T$:$L$] 44.1 131.4
(c) VLM layers[$K$:$K$+$T$] 57.4 153.6
Number of twig layers (Table 4d in main text)
(d) $T$ = 1 48.7 154.1
(e) $T$ = 2 53.4 152.6
(f) $T$ = 3 57.4 153.6
(g) $T$ = 4 58.1 145.4
Appendix B More Experimental Results
@@
More accuracy comparisons on benchmarks. To further demonstrate the superiority of TwigVLM, we present additional experimental results on a larger base VLM and more benchmarks, respectively.
Table 8 provides the accuracy comparisons on a larger LLaVA-1.5-13B model. TwigVLM consistently achieves the best overall RelAcc compared to all the counterparts, with its superiority being more significant as the increase of pruning ratios. These results verify the scalability and generalization ability of TwigVLM in accelerating large VLMs.
Taking LLaVA-1.5-7B as the base VLM, Table 9 compares the accuracies between TwigVLM and other visual token pruning methods on nine VLM benchmarks under three different pruning ratios. TwigVLM consistently outperforms or matches its counterparts on all benchmarks and pruning ratios, achieving the best overall RelAcc. In particular, TwigVLM even surpasses the upper bound given by the base VLM in RelAcc (100.6%) with a 66.7% pruning ratio, demonstrating its spectacular effectiveness and robustness in accelerating VLMs to deal with various tasks.
@@ Appendix D Evaluation Benchmarks
ActivityNet-QA [58] offers 58K human-annotated question-answer pairs across 5.8K videos from the ActivityNet dataset. Its questions include motion, spatial relationships, and temporal dynamics, requiring long-term spatio-temporal reasoning and making it an popular benchmark for evaluating advanced video understanding abilities. We denote it as “ActivityNet” in the main text.

### v2

@@ Abstract
Large vision-language models (VLMs) have demonstrated remarkable capabilities in open-world multimodal understanding, yet their high computational overheads pose great challenges for practical deployment. Some recent works have proposed methods to accelerate VLMs by pruning redundant visual tokens guided by the attention maps of VLM’s early layers. Despite the success of these token pruning methods, they still suffer from two major shortcomings: (i) considerable accuracy drop due to insensitive attention signals in early layers, and (ii) limited speedup when generating long responses (e.g., 30 tokens). To address the limitations above, we present TwigVLM—a simple and general architecture by “growing” a lightweight twig upon an early layer of the base VLM. Compared with most existing VLM acceleration methods purely based on visual token pruning, our TwigVLM not only achieves better accuracy retention by employing a twig-guided token pruning (TTP) strategy, but also yields higher generation speed by utilizing a self-speculative decoding (SSD) strategy. Taking LLaVA-1.5-7B as the base VLM, experimental results show that TwigVLM preserves 96% of the original performance after pruning 88.9% of visual tokens and achieves 154% speedup in generating long responses, delivering significantly better performance in terms of both accuracy and speed over the state-of-the-art VLM acceleration methods.
@@ 1 Introduction
Figure 1: Our TwigVLM overview. (a) Given a base deep VLM, the TwigVLM is obtained by freezing the base VLM while training a shallow twig block upon its early layer. (b) Compared to the VLM acceleration methods based on visual token pruning (e.g., FastV [9]), TwigVLM not only achieves better accuracy retention but also yields higher generation speed. Both FastV and TwigVLM take LLaVA-1.5-7B [32] as the base VLM (red dotted lines in the bar graphs) with the same pruning ratio of 88.9%.
The revolution in large language models (LLMs) has reshaped the landscape of artificial intelligence [4, 47, 18]. Meanwhile, there has been growing interest in building large vision-language models (VLMs) [41, 31, 1, 10, 56, 27, 44], which can perform visual understanding and reasoning and be used in various vision-language tasks such as object grounding [42, 38], document understanding [57, 37], GUI automation [50, 63], and robotic control [3, 24]. Despite the remarkable progress achieved by these large VLMs, they are usually parametric-intensive and computational-heavy. This poses great challenges for deploying these models in low-latency scenarios. Therefore, the research on VLM acceleration is of practical and emergent demands.
As the computational cost of transformer architectures [49] increases quadratically with token sequence length, reducing the number of tokens offers an effective way to accelerate VLMs. Moreover, visual tokens usually contain much more redundant information than textual tokens [9], especially in high-resolution images and long videos. Therefore, a series of works on visual token pruning has been proposed [9, 65, 34, 55]. For instance, FastV [9] leverages attention signals from an early layer of the VLM to drop redundant tokens for subsequent layers. Many following works have extended FastV’s idea to progressive token reduction across layers via pruning and merging operations [65, 34], while some others have been proposed to use the attention maps from the visual encoder to guide the visual token pruning before being input into the VLM backbone [55, 64]. Despite the success they have achieved, they still suffer from the following two major shortcomings:
@@
To this end, we present TwigVLM, a simple and general VLM acceleration approach by appending a lightweight twig block upon an early layer of the base VLM (see Figure 1(a)). After efficient post-training of the twig block only, the learned twig block is used in both the prefilling and decoding stages of inference. As shown in Figure 1(b), our TwigVLM not only achieves better accuracy retention by employing a twig-guided token pruning (TTP) strategy, but also yields higher generation speed by utilizing a self-speculative decoding (SSD) strategy [13]. To summarize, TwigVLM is easy to deploy and enables holistic inference acceleration with maximum accuracy retention.
@@ 4.3 Main Results
Results on LLaVA-1.5 and LLaVA-NeXT. As shown in Tables 1 and 2, we take LLaVA-1.5-7B and LLaVA-NeXT-7B as the base VLMs and compare TwigVLM with the existing VLM acceleration methods on six commonly-used VLM benchmarks. We can see that TwigVLM consistently and significantly outperforms all its counterparts under different pruning ratios. Notably, TwigVLM achieves near-perfect performance preservation (more than 99.0%) when pruning 66.7% visual tokens on LLaVA-1.5-7B and 77.8% tokens on LLaVA-NeXT-7B, showing the superiority of our TTP strategy over existing pruning strategies. Experiments on more benchmarks and VLMs (e.g., Qwen2.5-VL) are provided in the supplementary.
@@ 4.4 Ablation Studies
Number of twig layers. Table 4(d) varies the number of twig layers $T$ from 1 to 4 to investigate its effect on TwigVLM. We observe that a single twig layer ($T$=1) leads to the highest speed but the lowest accuracy. Increasing $T$ to 2 or 3 improves accuracy with negligible speed degradation. However, further increasing $T$ to 4 leads to a distinct performance drop in terms of both accuracy and speed. These observations can be explained by two factors: i) improving retained token selection (via more twig layers) obtains diminishing accuracy returns, which finally saturates at $T$=3, and ii) deeper twig block involves linearly increasing computation while the token acceptance ratio in SSD achieves saturation, leading to a decrease in generation speed. Thus, $T$=3 offers an optimal choice for the number of twig layers.
@@ 5 Related Work
Visual token reduction for VLMs. Unlike acceleration techniques for LLMs, token reduction methods for VLMs mainly address the problem of visual token redundancy to improve inference efficiency [68]. The pioneering work of FastV [9] first studies this redundancy phenomenon, revealing that pruning visual tokens at an early layer of the VLM can achieve acceleration while preserving the majority of the model’s accuracy. Following this paradigm, subsequent studies have introduced adaptive or hierarchical pruning strategies [65, 58, 34, 59, 51, 20] and training-based adoption [8, 53, 67, 6] to further reduce the accuracy drop. Meanwhile, there is another line of studies that performs more aggressive token reduction by pruning or merging visual tokens in the visual encoding stage [2, 43, 55, 64, 48]. However, the accuracy of these pruned models is suboptimal as they are unaware of the textual prompt when pruning visual tokens. Finally, although the above methods effectively accelerate the prefilling stage of VLMs, their speedup in the decoding stage is limited.
@@ 6 Conclusion
Acknowledgment
This work was supported in part by the National Natural Science Foundation of China under Grants (62422204, 62125201, U24B20174, 62402152, 62406093), the Key Research and Development Program of Zhejiang Province (No. 2025C01026), and in part by the Zhejiang Provincial Natural Science Foundation of China under Grant (LDT23F02025F02, LQN25F020017, LQ24F020032).
@@ Appendix A More Implementation Details
In our experiments, we use the LLaVA-665K dataset [32] to train the twig blocks for the LLaVA-1.5-7B and LLaVA-NeXT-7B models, and combine the LLaVA-665K and VideoInstruct-100K datasets [28, 39] for the Video-LLaVA-7B model. The optimization hyper-parameters are detailed in Table 5. All training is performed on a server equipped with 8 NVIDIA A100 GPUs. Under these conditions, the training is highly efficient, requiring only approximately 10% of the time needed to train the corresponding base VLM, e.g., training the twig block for the LLaVA-1.5-7B model takes about 10 GPU hours, while the training of the original LLaVA-1.5-7B takes about 100 GPU hours.
@@
[…]
Upper Bound, 1,280 Tokens  (100%)
Q2.5VL-7B 60.7 2347 82.7 75.3 83.2 77.9 100.0% 100.0%
[…]
FastV 57.2 2299 80.9 75.5 81.5 74.2 97.3% 101.7%
TwigVLM 60.4 2338 79.4 77.9 82.6 78.4 99.8% 147.7%
[…]
FastV 53.5 2246 78.6 75.3 79.2 70.6 94.1% 103.1%
TwigVLM 59.9 2238 77.4 78.1 81.4 78.4 98.3% 151.2%
[…]
FastV 45.1 1859 61.5 72.8 62.9 58.2 79.1% 104.3%
TwigVLM 57.6 2020 67.0 74.9 73.0 75.6 91.0% 152.3%
Table 9: Performance comparisons on Qwen2.5-VL-7B. The RelAcc and RelSpd are evaluated on the same benchmarks mentioned as the main text.
Appendix B More Experimental Results
B.1 More performance comparisons
Comparisons on more benchmarks. Taking LLaVA-1.5-7B as the base VLM, Table 8 compares the accuracies between TwigVLM and other visual token pruning methods on nine VLM benchmarks under three different pruning ratios. TwigVLM consistently outperforms or matches its counterparts on all benchmarks and pruning ratios, achieving the best overall RelAcc. In particular, TwigVLM even surpasses the upper bound given by the base VLM in RelAcc (100.6%) with a 66.7% pruning ratio, demonstrating its spectacular effectiveness and robustness in accelerating VLMs to deal with various tasks.
Comparisons on on more base VLMs. To further demonstrate the generalization ability and superiority of TwigVLM, we present additional experimental results on a larger VLM, LLaVA-1.5-13B, and a more recent VLM, Qwen2.5-VL-7B, which is very capable and representative.
Table 7 provides the accuracy comparisons on the larger LLaVA-1.5-13B model. TwigVLM consistently achieves the best overall RelAcc compared to all the counterparts, with its superiority being more significant as the increase of pruning ratios. These results verify the scalability and generalization ability of TwigVLM in accelerating large VLMs.
To evaluate the universality of TwigVLM to VLMs beyond the LLaVA family, we also apply it to Qwen2.5-VL-7B. Since the SFT data for QwenVL is large-scale yet inaccessible, we use a 5M subset of the MAmmoTH-VL-10M dataset (single image split) [19] as an alternative to train TwigVLM, which takes 12 hours on 8*A100 NVIDIA GPUs. The results in Table 9 show that TwigVLM can still maintain its prominent advantages over FastV in terms of accuracy and speed, showing the effectiveness of TwigVLM in practical scenarios.
B.2 More ablation studies
ablation variant   TokAR (%) RelSpd (%)
Twig block initialization (Table 4c in main text)
(a) random init. 37.7 120.4
(b) VLM layers[$L$-$T$:$L$] 44.1 131.4
(c) VLM layers[$K$:$K$+$T$] 57.4 153.6
Number of twig layers (Table 4d in main text)
(d) $T$ = 1 48.7 154.1
(e) $T$ = 2 53.4 152.6
(f) $T$ = 3 57.4 153.6
(g) $T$ = 4 58.1 145.4
Data efficiency. To demonstrate the data efficiency of TwigVLM, we train multiple models using different proportions (i.e., 25%, 50%, 75%, and 100%) of each model’s respective training dataset for LLaVA-1.5-7B and Qwen2.5-VL-7B. As shown in Figure 7, both models exhibit a general upward trend in accuracy and speed as the amount of training data increases. Remarkably, however, even when trained on only 50% of their respective datasets, TwigVLM models already achieve competitive, and in some cases comparable, performance to models trained on the full dataset. Moreover, TwigVLM requires only 10% of the training cost of the corresponding base VLM (see A). To sum up, it is highly efficient and feasible to apply TwigVLM in industrial scenarios.
(a) TwigVLM on LLaVA-1.5-7B
(b) TwigVLM on Qwen2.5-VL-7B
Figure 7: Performance comparisons of TwigVLM models trained with different proportions of the training dataset. Specifically, we use LLaVA-665K to train TwigVLM models for LLaVA-1.5-7B and use MAmmoTH-VL-10M to train TwigVLM models for Qwen2.5-VL-7B. Even with only 50% of respective training data, TwigVLM is able to maintain competitive accuracy and speed.
Memory footprint analysis. We measure the inference VRAM usage of the LLaVA-1.5-7B and LLaVA-Next-7B models in Table 11. The introduction of the twig block brings 8% extra VRAM cost for loading model weights. Compared to the base VLM, the overall inference VRAM cost of TwigVLM is comparable or slightly reduced due to the substantial reduction of visual tokens.
model avg. visual tokens ($\bar{R}$) model weights VRAM (GB) inference VRAM (GB)
LLaVA-1.5-7B 576 14.3 15.8
+ TwigVLM 64 15.5 16.5
LLaVA-Next-7B 2,880 14.3 17.9
+ TwigVLM 320 15.5 16.8
Table 11: Memory footprint comparisons during inference.
@@ Appendix D Evaluation Benchmarks
ActivityNet-QA. ActivityNet-QA [61] offers 58K human-annotated question-answer pairs across 5.8K videos from the ActivityNet dataset. Its questions include motion, spatial relationships, and temporal dynamics, requiring long-term spatio-temporal reasoning and making it an popular benchmark for evaluating advanced video understanding capabilities. We denote it as “ActivityNet” in the main text.